### v1

@@ 6.1.1 Overview
We now show that the first result of Theorem 6.2 holds when we stop recurring on a sub-cluster. Assume that when running in some sub-cluster $R\subseteq S$, Corollary 6.5 terminated with case (1). In this case, $G_{\tilde{R}}$ is $\Omega(\phi_{R})$-expander with respect to $\mu_{R}$, which means that $X_{B_{R}}$ $\Omega(\phi_{R})$-expands in $G_{\tilde{R}}$. Therefore, by Remark 3.6, $G_{\tilde{R}}$ $\Omega(\phi_{R})$-respects the all-to-all flow problem on the vertices of $X_{B_{R}}$, which establishes the first result. Now, assume that Corollary 6.5 terminated with case (3b). This means that we have $A\subseteq R$ on which we do not recur, which satisfies that $G_{\tilde{A}}$ is an $\Omega(\phi_{R})$-expander with respect to $\mu_{R}$. Equivalently, $\tilde{A}\cap X_{B_{R}}$ $\Omega(\phi_{R})$-expands in $G_{\tilde{A}}$. Denote by $C=E(A,R\setminus A)$ the set of cut edges that were cut in this step. We have that we can route a flow in $G_{\tilde{A}}$ with congestion $O(1)$, such that each edge $e\in C$ sends one unit of flow, and each $v\in\tilde{A}$ gets at most $O(\phi_{R}\cdot\mu_{R}(v))=O(\mu_{R}(v))$ units.  [^22] Recall from Section 3.3 the definition of routing a flow from the cut edges. This means that only boundary nodes in $\tilde{A}\cap X_{B_{R}}$ can receive flow, and each receives at most $O(1)$ units. Let $X_{C}$ denote the new boundary edges in $G_{\tilde{A}}$, that correspond to the cut edges of $C$. Using Lemma 4.3, we get that $(\tilde{A}\cap X_{B_{R}})\cup X_{C}=X_{B_{A}}$ $\Omega(\phi_{R})$-expands in $G_{\tilde{A}}$. This means (Remark 3.6) that $G_{\tilde{A}}$ $\Omega(\phi_{R})$-respects the all-to-all flow problem on the vertices of $X_{B_{A}}$. Finally, note that because $|A|\leq|R|$, we have $\phi_{R}=\Omega(\phi_{A})$.
Next, we show the second result of Theorem 6.2. The argument follows the ideas of [harrelson2003polynomial, RackeS14]. Note that inter-cluster edges are only created as cut edges in cases (2a), (2b), (2c), (3a) and (3b) of Corollary 6.5.
We view the cuts formed by cutting $S$ as a binary tree [^23] This tree is for analysis purposes and should not be confused with the outer hierarchical decomposition., denoted the refinement tree $T_{S}^{\mathrm{ref}}$ of $S$ (see Figure 8). The root of the binary tree corresponds to $S$ itself and every vertex corresponds to a subset $D\subseteq S$. The children of $D$ in the binary tree are the subsets we recur on during the application of Corollary 6.5 on $G_{\tilde{D}}$. Note that in case (2c) we have that $D$ splits into $A_{1},D\setminus A_{1}$, and $A_{1}$ further splits into $A_{2},A_{1}\setminus A_{1}$. These are two levels of $T_{S}^{\mathrm{ref}}$. We choose to set the left child of each node to be the side in which we can route flow from the cut edges:
@@
Note that in all cases the left child $L\subseteq D$ satisfies $|L|\leq\frac{3}{4}|D|$, and additionally, we can route in $G_{\tilde{L}}$ a flow with congestion $O(1)$ such that each cut edge of $E(L,D\setminus L)$ sends one unit of flow and each vertex $v\in\tilde{L}$ receives at most $O(\phi_{D}\log n\cdot\mu_{D}(v))$ units. [^24] In case (2c), this is true for $D$ directly, while for $A_{1}$ it’s true with $\phi_{D}$ and $\mu_{D}$ instead of $\phi_{A_{1}}$ and $\mu_{A_{1}}$. However, because $|A_{1}|=\Theta(|D|)$, we have $\phi_{A_{1}}=\Theta(\phi_{D})$. Additionally, we have $\mu_{D}\leq\mu_{A_{1}}$. That is, each boundary split node of $X_{B_{D}}\cap\tilde{L}$ gets at most $O(\phi_{D}\log n)$ units. Let $c_{0}\geq 1$ be such that in all cases each boundary split node in $X_{B_{D}}\cap\tilde{L}$ gets at most $c_{0}\phi_{D}\log n$ units. In the definition of $f$ in Theorem 6.2, we will choose $c_{f}\mathrel{\mathop{:}}=\frac{4c_{0}}{\log(\frac{4}{3})}\geq 1$.
@@
Note that the left depth of a node is at most $\log_{4/3}{n}\leq 3\log n$. Therefore, any node of depth $3\log n$ must be a leaf node, so $\Phi(3\log n)=1$ (no routings are done at the leaf level). Following the analysis of [harrelson2003polynomial], we now prove by induction on $j=3\log n,\ldots,1$ the following claim:
Claim 6.6 (See [harrelson2003polynomial, Lemma 4]).
@@ A.1 Omitted Proofs from Section 3
We now prove that $\operatorname{dem}_{P^{\prime}}(B,W)\leq\operatorname{dem}_{Q}(B,W)$, for every cut $(B,W)$. Fix a cut $(B,W)$, we get that:
$\operatorname{dem}_{P^{\prime}}(B,W)$ $=\sum_{k\in\mathcal{K}}\left|\sum_{u\in B}P^{\prime}(u,k)\right|$
$=\sum_{k\in\mathcal{K}}\left|\sum_{u\in B}\sum_{v\in V}{\frac{P(v,k)}{\left\|P(v)\right\|_{1}}D_{Q}(v,u)}-\sum_{u\in B}\sum_{v\in V}{\frac{P(u,k)}{\left\|P(u)\right\|_{1}}D_{Q}(u,v)}\right|$
$=\sum_{k\in\mathcal{K}}\left|\sum_{u\in B}\sum_{v\in B}{\frac{P(v,k)}{\left\|P(v)\right\|_{1}}D_{Q}(v,u)}+\sum_{u\in B}\sum_{v\in W}{\frac{P(v,k)}{\left\|P(v)\right\|_{1}}D_{Q}(v,u)}\right$.
$\;\;\;\;\;\;\;\left.-\sum_{u\in B}\sum_{v\in B}{\frac{P(u,k)}{\left\|P(u)\right\|_{1}}D_{Q}(u,v)}-\sum_{u\in B}\sum_{v\in W}{\frac{P(u,k)}{\left\|P(u)\right\|_{1}}D_{Q}(u,v)}\right|$
$=\sum_{k\in\mathcal{K}}\left|\sum_{u\in B}\sum_{v\in W}{\frac{P(v,k)}{\left\|P(u)\right\|_{1}}D_{Q}(v,u)}-\sum_{u\in B}\sum_{v\in W}{\frac{P(u,k)}{\left\|P(u)\right\|_{1}}D_{Q}(u,v)}\right|$
$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu1\mkern-1.5mu)}}{{\leq}}\sum_{u\in B}\sum_{v\in W}\sum_{k\in\mathcal{K}}{\frac{|P(v,k)|}{\left\|P(v)\right\|_{1}}D_{Q}(v,u)+\frac{|P(u,k)|}{\left\|P(u)\right\|_{1}}D_{Q}(u,v)}$
$=\sum_{u\in B}\sum_{v\in W}\sum_{k\in\mathcal{K}}{D_{Q}(v,u)+D_{Q}(u,v)}=\operatorname{dem}_{Q}(B,W),$
@@
A.2 Omitted Proofs from Section 5
@@
Recall Figure 3. Intuitively, when we spread the demands of $P(v)$ on the neighboring split nodes of $v$, we do not change the demand across the $(B^{\prime},W^{\prime})$ cut, except in the case that this spreading moved mass from one side of the cut to the other. In other words, this happens only if the edge $(x_{e},v)$ crosses the cut. This may occur only if $e=(u,v)\in E(B,W)$, and $v\in W$. In this case, the amount of mass that crosses from $v$ to $x_{e}$ is at most $1$. Summing across all edges in $E(B,W)$, we only change the demand across the cut by at most $\operatorname{cap}(B,W)$.
@@
$\operatorname{dem}_{P_{\mathcal{Q}_{0}}}(B^{\prime},W^{\prime})$ $=\sum_{k\in\mathcal{K}}\left|\sum_{u\in B^{\prime}}P_{\mathcal{Q}_{0}}(u,k)\right|=\sum_{k\in\mathcal{K}}\left|\sum_{u\in B^{\prime}}\sum_{v\in V}P_{\{v\}^{\prime}}(u,k)\right|$
$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu1\mkern-1.5mu)}}{{\geq}}\sum_{k\in\mathcal{K}}\left|\sum_{v\in B}\sum_{u\in B^{\prime}}P_{\{v\}^{\prime}}(u,k)\right|-\sum_{k\in\mathcal{K}}\left|\sum_{v\in W}\sum_{u\in B^{\prime}}P_{\{v\}^{\prime}}(u,k)\right|$
$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu2\mkern-1.5mu)}}{{\geq}}\sum_{k\in\mathcal{K}}\left|\sum_{v\in B}P(v,k)\right|-\sum_{k\in\mathcal{K}}\sum_{v\in W}\sum_{u\in B^{\prime}}\left|P_{\{v\}^{\prime}}(u,k)\right|$
$=\operatorname{dem}_{P}(B,W)-\sum_{v\in W}\sum_{k\in\mathcal{K}}\sum_{u\in B^{\prime}\cap\{v\}^{\prime}}\left|P_{\{v\}^{\prime}}(u,k)\right|$
$=\operatorname{dem}_{P}(B,W)-\sum_{v\in W}\sum_{u\in B^{\prime}\cap\{v\}^{\prime}}\sum_{k\in\mathcal{K}}\left|P_{\{v\}^{\prime}}(u,k)\right|$
$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu3\mkern-1.5mu)}}{{\geq}}\operatorname{dem}_{P}(B,W)-\sum_{v\in W}\sum_{u\in B^{\prime}\cap\{v\}^{\prime}}1$
$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu4\mkern-1.5mu)}}{{=}}\operatorname{dem}_{P}(B,W)-\operatorname{cap}_{G}(B,W),$
where Inequality $(1)$ holds by swapping the order of summation and by triangle inequality; Inequality $(2)$ holds by the way we defined $P_{\{v\}^{\prime}}$ (spreading out the demands of $P(v)$ on the adjacent split nodes), as well as the way we defined $B^{\prime}$ (that is, $B^{\prime}$ contains all the split nodes defined by the cut $(B,W)$ in $G$) and triangle inequality; Inequality $(3)$ holds by the mentioned observation that $\left\|P_{\{v\}^{\prime}}(x_{e})\right\|_{1}\leq 1$, and Equality $(4)$ holds by the way we defined $B^{\prime}$.
A.3 Algebraic Tools
Claim A.1 (See [harrelson2003polynomial, Lemma 10]).

### v2

@@ 6.1.1 Overview
We now show that the first claim of Theorem 6.2 holds when we stop recurring on a sub-cluster. Assume that when running in some sub-cluster $R\subseteq S$, Corollary 6.5 terminated with Case (1). In this case, $G_{\tilde{R}}$ is $\Omega(\phi_{R})$-expander with respect to $\mu_{R}$, which means that $X_{B_{R}}$ $\Omega(\phi_{R})$-expands in $G_{\tilde{R}}$. Therefore, by Remark 3.6, $G_{\tilde{R}}$ $\Omega(\phi_{R})$-respects the all-to-all flow problem on the vertices of $X_{B_{R}}$, which establishes the first result.
Now, assume that Corollary 6.5 terminated with case (3b). This means that we have $A\subseteq R$ on which we do not recur, which satisfies that $G_{\tilde{A}}$ is an $\Omega(\phi_{R})$-expander with respect to $\mu_{R}$. Equivalently, $\tilde{A}\cap X_{B_{R}}$ $\Omega(\phi_{R})$-expands in $G_{\tilde{A}}$. Denote by $C=E(A,R\setminus A)$ the set of cut edges that were cut in this step. We have that we can route a flow in $G_{\tilde{A}}$ with congestion $O(1)$, such that each edge $e\in C$ sends one unit of flow, and each $v\in\tilde{A}$ gets at most $O(\phi_{R}\cdot\mu_{R}(v))=O(\mu_{R}(v))$ units. [^29] Recall the definition of routing a flow from the cut edges, introduced in the discussion following Theorem 3.9. This means that only boundary nodes in $\tilde{A}\cap X_{B_{R}}$ can receive flow, and each receives at most $O(1)$ units. Let $X_{C}$ denote the new boundary edges in $G_{\tilde{A}}$, that correspond to the cut edges of $C$. Using Lemma 4.3, we get that $(\tilde{A}\cap X_{B_{R}})\cup X_{C}=X_{B_{A}}$ $\Omega(\phi_{R})$-expands in $G_{\tilde{A}}$. This means (Remark 3.6) that $G_{\tilde{A}}$ $\Omega(\phi_{R})$-respects the all-to-all flow problem on the vertices of $X_{B_{A}}$. Finally, since $|A|\leq|R|$, we have $\phi_{R}=\Omega(\phi_{A})$. That is, $G_{\tilde{A}}$, $\Omega(\phi_{A})$-respects the all-to-all flow problem on the vertices of $X_{B_{A}}$, as desired.
Next, we show the second claim of Theorem 6.2. The argument follows the ideas of [15, 28]. Note that inter-cluster edges are only created as cut edges in cases (2a), (2b), (2c), (3a) and (3b) of Corollary 6.5.
We view the cuts formed by cutting $S$ as a binary tree. [^30] This tree is for analysis purposes and should not be confused with the outer hierarchical decomposition. We call it the refinement tree, $T_{S}^{\mathrm{ref}}$, of $S$ (see Figure 8). The root of the binary tree corresponds to $S$ itself and every vertex corresponds to a subset $D\subseteq S$. The children of $D$ in the binary tree are the subsets we recur on during the application of Corollary 6.5 on $G_{\tilde{D}}$. Note that in case (2c) we have that $D$ splits into $A_{1},D\setminus A_{1}$, and $A_{1}$ further splits into $A_{2},A_{1}\setminus A_{1}$. These correspond to two levels of $T_{S}^{\mathrm{ref}}$. We choose to set the left child of each node to be the side in which we can route flow from the cut edges:
@@
Note that in all cases the left child $L\subseteq D$ satisfies $|L|\leq\frac{3}{4}|D|$, and additionally, we can route in $G_{\tilde{L}}$ a flow with congestion $O(1)$ such that each cut edge of $E(L,D\setminus L)$ sends one unit of flow and each vertex $v\in\tilde{L}$ receives at most $O(\phi_{D}\log n\cdot\mu_{D}(v))$ units. [^31] In case (2c), this is true for $D$ directly, while for $A_{1}$ it’s true with $\phi_{D}$ and $\mu_{D}$ instead of $\phi_{A_{1}}$ and $\mu_{A_{1}}$. However, because $|A_{1}|=\Theta(|D|)$, we have $\phi_{A_{1}}=\Theta(\phi_{D})$. Additionally, we have $\mu_{D}\leq\mu_{A_{1}}$. That is, each boundary split node of $X_{B_{D}}\cap\tilde{L}$ gets at most $O(\phi_{D}\log n)$ units. Let $c_{0}\geq 1$ be such that in all cases each boundary split node in $X_{B_{D}}\cap\tilde{L}$ gets at most $c_{0}\phi_{D}\log n$ units. In the definition of $f$ in Theorem 6.2, we choose $c_{f}\mathrel{\mathop{:}}=\frac{4c_{0}}{\log(\frac{4}{3})}\geq 1$.
@@
Note that the left depth of a node is at most $\log_{4/3}{n}\leq 3\log n$. Therefore, any node of depth $3\log n$ must be a leaf node, so $\Phi(3\log n)=1$ (no routings are done at the leaf level). Following the analysis of [15], we now prove by induction on $j=3\log n,\ldots,1$ the following claim:
Claim 6.6 (See [15, Lemma 4]).
@@ A.1 Omitted Proofs from Section 2
To prove Claim 2.11, we use the following flow-cut gap theorems [19, 4, 24]:
Theorem A.1 (Flow-Cut Gap [4]).
Let $Q$ be a demand matrix on $G=(V,E)$. Define $\theta_{G}(Q)\ :=\ \sup\{\theta:\text{$G$ $\theta$-respects $Q$}\}.$ Then, for a universal constant $c>0$,
Theorem A.2 (Flow-Cut Gap of Trees [24]).
Let $Q$ be a demand matrix on a tree $T=(V,E)$. Define $\theta_{T}(Q)\ :=\ \sup\{\theta:\text{$T$ $\theta$-respects $Q$}\}.$ Then, $\frac{1}{\theta_{T}(Q)}=\operatorname{cong}_{T}(Q)$.
Proof of Claim 2.11.
We begin by proving the first item. Let $T$ be a tree cut-sparsifier of $G$ of quality $\alpha$. Using Theorem A.1, we get that a tree cut-sparsifier $T$ of $G$ of quality $\alpha$ satisfies, for all flow problems $Q$,
$\operatorname{cong}_{T}(Q)\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu1\mkern-1.5mu)}}{{\leq}}\operatorname{cong}_{G}(Q)\leq\frac{\log n}{c\cdot\theta_{G}(Q)}\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu2\mkern-1.5mu)}}{{\leq}}\frac{\alpha\log n}{c\cdot\theta_{T}(Q)}\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu3\mkern-1.5mu)}}{{=}}\frac{\alpha\log n}{c}\operatorname{cong}_{T}(Q),$
where Inequality (2) holds since $T$ is a cut-sparsifier of quality $\alpha$, and therefore
$\frac{\operatorname{mincut}_{T}(S,\bar{S})}{\operatorname{dem}_{Q}(S,\bar{S})}\leq\frac{\alpha\cdot\operatorname{cap}_{G}(S,\bar{S})}{\operatorname{dem}_{Q}(S,\bar{S})},$
for every cut $S\subseteq V$ with $\operatorname{dem}_{Q}(S,\bar{S})>0$. Equality (3) holds by Theorem A.2. Finally, Inequality (1) holds since, for every $e=\{S,P_{T}(S)\}\in T$, the congestion on $e$ is exactly $\frac{\operatorname{cap}_{G}(S,V\setminus S)}{\operatorname{dem}_{Q}(S,V\setminus S)}$, which trivially lower bounds $\operatorname{cong}_{G}(Q)$. Therefore, $T$ is a flow-sparsifier of $G$ of quality $O(\alpha\log n)$.
We proceed to prove the second item. Let $T$ be a tree flow-sparsifier of $G$ of quality $\alpha$ and fix a cut $(S,\bar{S})$ of $V$. We prove that $\operatorname{cap}_{G}(S,\bar{S})\leq\operatorname{mincut}_{T}(S,\bar{S})\leq\alpha\cdot\operatorname{cap}_{G}(S,\bar{S})$. The first inequality was already established in the proof of Claim 5.2 and relies only on the definition of the capacities in $T$. For the second inequality, we reuse the same demand matrix $R$ constructed in the proof of Claim 5.2. Namely, $R$ is the demand matrix induced by a maximum $(s,t)$-flow in the auxiliary network $T^{\prime}=T\cup\{s,t\}$, where $s$ is connected to all vertices in $\bar{S}$ and $t$ to all vertices in $S$ with infinite-capacity edges. By construction, $\operatorname{dem}_{R}(S,\bar{S})=\operatorname{mincut}_{T}(S,\bar{S})$, and $\operatorname{cong}_{T}(R)=1$. Since $T$ is a flow-sparsifier of $G$ of quality $\alpha$, we get $\operatorname{cong}_{G}(R)\leq\alpha$. Therefore
$\frac{\operatorname{mincut}_{T}(S,\bar{S})}{\operatorname{cap}_{G}(S,\bar{S})}=\frac{\operatorname{dem}_{R}(S,\bar{S})}{\operatorname{cap}_{G}(S,\bar{S})}\leq\operatorname{cong}_{G}(R)=\alpha.$
A.2 Omitted Proofs from Section 3
@@
We now prove that $\operatorname{dem}_{P^{\prime}}(U,W)\leq\operatorname{dem}_{Q}(U,W)$, for every cut $(U,W)$. Fix a cut $(U,W)$, we get that:
$\operatorname{dem}_{P^{\prime}}(U,W)$ $=\sum_{k\in\mathcal{K}}\left|\sum_{u\in U}P^{\prime}(u,k)\right|$
$=\sum_{k\in\mathcal{K}}\left|\sum_{u\in U}\sum_{v\in V}{\frac{P(v,k)}{\left\|P(v)\right\|_{1}}D_{Q}(v,u)}-\sum_{u\in U}\sum_{v\in V}{\frac{P(u,k)}{\left\|P(u)\right\|_{1}}D_{Q}(u,v)}\right|$
$=\sum_{k\in\mathcal{K}}\left|\sum_{u\in U}\sum_{v\in U}{\frac{P(v,k)}{\left\|P(v)\right\|_{1}}D_{Q}(v,u)}+\sum_{u\in U}\sum_{v\in W}{\frac{P(v,k)}{\left\|P(v)\right\|_{1}}D_{Q}(v,u)}\right$.
$\;\;\;\;\;\;\;\left.-\sum_{u\in U}\sum_{v\in U}{\frac{P(u,k)}{\left\|P(u)\right\|_{1}}D_{Q}(u,v)}-\sum_{u\in U}\sum_{v\in W}{\frac{P(u,k)}{\left\|P(u)\right\|_{1}}D_{Q}(u,v)}\right|$
$=\sum_{k\in\mathcal{K}}\left|\sum_{u\in U}\sum_{v\in W}{\frac{P(v,k)}{\left\|P(u)\right\|_{1}}D_{Q}(v,u)}-\sum_{u\in U}\sum_{v\in W}{\frac{P(u,k)}{\left\|P(u)\right\|_{1}}D_{Q}(u,v)}\right|$
$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu1\mkern-1.5mu)}}{{\leq}}\sum_{u\in U}\sum_{v\in W}\sum_{k\in\mathcal{K}}{\frac{|P(v,k)|}{\left\|P(v)\right\|_{1}}D_{Q}(v,u)+\frac{|P(u,k)|}{\left\|P(u)\right\|_{1}}D_{Q}(u,v)}$
$=\sum_{u\in U}\sum_{v\in W}{D_{Q}(v,u)+D_{Q}(u,v)}=\operatorname{dem}_{Q}(U,W),$
@@
A.3 Omitted Proofs from Section 5
@@
Recall Figure 3. Intuitively, when we spread the demands of $P(v)$ on the neighboring split nodes of $v$, we do not change the demand across the $(U^{\prime},W^{\prime})$ cut, except in the case that this spreading moved mass from one side of the cut to the other. In other words, this happens only if the edge $(x_{e},v)$ crosses the cut. This may occur only if $e=(u,v)\in E(U,W)$, and $v\in W$. In this case, the amount of mass that crosses from $v$ to $x_{e}$ is at most $1$. Summing across all edges in $E(U,W)$, we only change the demand across the cut by at most $\operatorname{cap}(U,W)$.
@@
$\operatorname{dem}_{P_{\mathcal{Q}_{0}}}(U^{\prime},W^{\prime})$ $=\sum_{k\in\mathcal{K}}\left|\sum_{u\in U^{\prime}}P_{\mathcal{Q}_{0}}(u,k)\right|=\sum_{k\in\mathcal{K}}\left|\sum_{u\in U^{\prime}}\sum_{v\in V}P_{\{v\}^{\prime}}(u,k)\right|$
$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu1\mkern-1.5mu)}}{{\geq}}\sum_{k\in\mathcal{K}}\left|\sum_{v\in U}\sum_{u\in U^{\prime}}P_{\{v\}^{\prime}}(u,k)\right|-\sum_{k\in\mathcal{K}}\left|\sum_{v\in W}\sum_{u\in U^{\prime}}P_{\{v\}^{\prime}}(u,k)\right|$
$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu2\mkern-1.5mu)}}{{\geq}}\sum_{k\in\mathcal{K}}\left|\sum_{v\in U}P(v,k)\right|-\sum_{k\in\mathcal{K}}\sum_{v\in W}\sum_{u\in U^{\prime}}\left|P_{\{v\}^{\prime}}(u,k)\right|$
$=\operatorname{dem}_{P}(U,W)-\sum_{v\in W}\sum_{k\in\mathcal{K}}\sum_{u\in U^{\prime}\cap\{v\}^{\prime}}\left|P_{\{v\}^{\prime}}(u,k)\right|$
$=\operatorname{dem}_{P}(U,W)-\sum_{v\in W}\sum_{u\in U^{\prime}\cap\{v\}^{\prime}}\sum_{k\in\mathcal{K}}\left|P_{\{v\}^{\prime}}(u,k)\right|$
$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu3\mkern-1.5mu)}}{{\geq}}\operatorname{dem}_{P}(U,W)-\sum_{v\in W}\sum_{u\in U^{\prime}\cap\{v\}^{\prime}}1$
$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu4\mkern-1.5mu)}}{{=}}\operatorname{dem}_{P}(U,W)-\operatorname{cap}_{G}(U,W),$
where Inequality $(1)$ holds by swapping the order of summation and by triangle inequality; Inequality $(2)$ holds by the way we defined $P_{\{v\}^{\prime}}$ (spreading out the demands of $P(v)$ on the adjacent split nodes), as well as the way we defined $U^{\prime}$ (that is, $U^{\prime}$ contains all the split nodes defined by the cut $(U,W)$ in $G$) and triangle inequality; Inequality $(3)$ holds by the mentioned observation that $\left\|P_{\{v\}^{\prime}}(x_{e})\right\|_{1}\leq 1$, and Equality $(4)$ holds by the way we defined $U^{\prime}$.
A.4 Algebraic Tools
Claim A.3 (See [15, Lemma 10]).